# Cocoa: Co-Planning and Co-Execution with AI Agents

### v3

@@ 2.3. AI for Scientific Research
Significant efforts in recent years have advanced how AI can be used to support scientific research. Of particular interest to us are works that leverage these recent advances to help researchers with planning and execution of literature-focused tasks. Planning a research project is a cognitively demanding, complex process consisting of iterative cycles of divergent and convergent thinking grounded in literature review (Palani et al., 2023; Dean et al., 2006; Chang et al., 2023), making it an ideal scenario to explore human-AI collaboration.
@@
One major thread of prior research has focused on leveraging AI models as tools or components in user-driven systems. These include tools for paper reading (Chang et al., 2023; Rachatasumrit et al., 2022) and skimming (Fok et al., 2022), literature review (Lee et al., 2024b; Palani et al., 2023), paper recommendation (Kang et al., 2022b, a), information retrieval and sensemaking (Chan et al., 2018; Kang et al., 2023; Fok et al., 2023; Wang et al., 2024a), ideation (Gu and Krenn, 2024; Baek et al., 2024; Jansen et al., 2024) and “deep research” systems that combine support for multiple activities (Google, 2025; OpenAI, 2025a; Perplexity, 2025). Progress in this area has given rise to excitement for a “computational inflection for scientific discovery” (Hope et al., 2022). On the other end of the spectrum are fully automated end-to-end systems where an AI agent has much more agency in planning and executing tasks to attempt to carry out entire research projects on their own (Lu et al., 2024; Jansen et al., 2025; Gottweis et al., 2025).
@@
As AI systems become more capable, we ask this question: Is the generation of full research artifacts (e.g., research questions, proposals) the most desired and promising paradigm to assist researchers? Indeed, when working with other (human) collaborators, researchers often find richness in co-evolving partially completed artifacts (Star and Griesemer, 1989; Lee, 2007; Shrum et al., 2007; Jirotka et al., 2013) and answering or asking questions that stimulate critical thinking and reflection (Overholser, 1993; Padesky, 1993). Further, some fully automatically generated artifacts are still found to be of lower quality than those generated with human involvement (Jansen et al., 2025). A limited number of works have thus explored human-AI co-creation of research artifacts (Babaian et al., 2002; Liu et al., 2024a; Nigam et al., 2024; Kang et al., 2022b). These include in-document commands that trigger an AI assistant to complete a partial citation based on the user’s preferred bibliographies and paper collections (Babaian et al., 2002), interactive node-based editors to iteratively expand upon and refine AI-generated research questions (Liu et al., 2024a; Pu et al., 2024), and emulating colleague and mentor personas using LLMs to work with the user in developing research proposals (Nigam et al., 2024).
@@
However, despite these systems, there is little clarity on how researchers want to co-create with AI during research planning and execution. Prior work (e.g., (Liu et al., 2024a; Nigam et al., 2024)) assumed that generating particular artifacts is helpful to researchers. Rather than focusing on generating predetermined research artifacts, we contribute an interaction design pattern that allows for flexible specification of the final artifact and collaboration between a researcher and an AI agent.
@@ 3. Formative Study
Building on prior work, we aim to enhance collaboration between humans and AI agents using a shared, interactive operational representation. We focus on scientific research—a challenging domain for planning given its complex, multi-step workflows and the nuanced decisions researchers make upon processing information-dense scientific papers. Specifically, we investigate the potential for project documents—continuous records of ideas, updates, and tasks—to act as the agent environment. We thus conducted a formative study to answer the following research questions:
@@ 3.1. Participants, Procedure, and Data Analysis
We recruited 9 Ph.D. students (detailed demographics in Table 1 of Appendix B) through an interest form sent to a research organization’s internal Slack channel and the authors’ professional connections. We targeted Ph.D. students as they often lead the detailed planning and execution of research projects and are the primary editors of project documents.
@@
Details for each part can be found in Appendix C. Each participant was given a \$35 USD honorarium after the study. The study was reviewed and exempted by our organization’s internal review board. We qualitatively analyzed project documents produced from the study probe activity alongside study transcripts. Our data analysis methods can be found in Appendix D.
@@ 3.2. Findings
3.2.1. [F1 & F3] Project documents are promising environments for human-agent interaction
@@ 3.3. Design Goals
  • DG2:
    Allow flexible delegation of agency between researcher and agent. Researchers may not want to delegate all parts of a plan to AI [F2, F3]. This preference may also be context-dependent and constantly shifting. Our system uses a concrete implementation of a flexible delegation approach proposed by Satyanarayan (Satyanarayan, 2024).
  • DG3:
    Provide opportunities for researcher-AI collaboration in both planning and execution. Agent developers have called for decoupling planning and execution to avoid fruitless execution (Huyen, 2025). Indeed, participants discussed planning (i.e., breaking down a problem into smaller, more actionable tasks like they did in their project documents) and execution (i.e., completing steps in the plan and refining outputs, with or without AI assistance) as separate but complementary actions. We thus see planning and execution as two distinct stages for fertile researcher-AI collaboration. Because these two stages are closely intertwined and synergistic, we aim to interleave the two in our system.
@@ 4. Cocoa: System Walkthrough and Implementation
We present Cocoa, a system that embeds an AI agent into a document editor using a novel interaction design pattern we call interactive plans. Interactive plans allow users to collaboratively plan (co-plan) with the agent—the agent proposes an initial plan of execution to tackle a user request that the user can edit to their liking. Then, users can collaboratively execute (co-execute) the plan with the agent—the user and the agent can build off each other’s work to synergize human and AI capabilities, and provide flexible negotiation of human and machine agency.
To illustrate the features and functionality of Cocoa, we follow the journey of Nóírín, a researcher in human-AI interaction, as she uses her project document to further explore open questions in interactive interfaces for AI agents.
@@ 4.1. Co-Planning
4.1.1. Invoking the agent and selecting plans
Nóírín’s project document is an informal, reverse chronological log of research progress and includes information such as meeting notes, rough ideas, links to literature, and questions for herself and her collaborators. She has a specific question in her document that she wants to explore further, but is not sure how to proceed on her own: “What are new ways AI agents can interactively elicit human feedback?”
To initiate co-planning, she highlights the question to invoke the agent on it using a button that appears in a floating menu. The agent analyzes the question within the context of her entire document and proposes a few different plans for approaching it. These plans are displayed in the document via a plan selector UI within the editor, which allows Nóírín to browse and select a plan. Upon Nóírín’s selection, a plan becomes fully interactive within the document (Figure 3).
@@ 4.1.1. Invoking the agent and selecting plans
4.1.2. Agent steps and user steps
The initial plan included steps for brainstorming relevant topics, searching for papers, and making connections between papers (e.g., Figure 3). Nóírín wants to assign tasks that are relatively low risk but high effort to the agent, such as searching for papers and expanding on some of her preliminary ideas, and keep tasks requiring higher-level thinking and deep research expertise to herself, such as identifying papers that can seed deeper exploration and surfacing connections between papers. To do so, she uses the step assignment toggle (Figure 4) to assign the step to herself (a “user step”) or the agent (an “agent step”). User steps are highlighted in blue. When executing the plan, the agent will automatically attempt agent steps but will request Nóírín’s input on user steps.
Figure 4. For each plan step, the user can assign the step to either themselves (a user step) or the agent (an agent step) by using the step assignment toggle, as well as edit the step’s description.
[…]
An agent step (plan step assigned to the agent) is shown above a user step (plan step assigned to the user). The user can assign the step to either themselves or the agent by using the step assignment toggle beside the step, as well as edit the step’s description directly, just like they were editing text elsewhere in the document.
4.1.3. Editing the plan step description
Upon further review of the plan, Nóírín is now curious about authors who have published on “interactive feedback mechanisms in AI” rather than seeing papers on the topic directly. She edits the step description to reflect this. The step description can be edited just like editing any other bulleted list in her document, making this intuitive for Nóírín. If she chooses, Nóírín can also employ LLM assistance to suggest a step to replace the current one, informed by the previous steps (Figure 5). Nóírín can easily undo the suggestion if she prefers the original step. Once Nóírín is satisfied with her editing, she saves her changes with the keyboard shortcut Cmd/Ctrl+S to register them with Cocoa.
Figure 5. Highlighting text within the step description will bring up an option for the agent to suggest an alternate step. The user can undo the suggestion or save to accept it.
[…]
The user highlights text within a plan step, revealing a button labelled ’Suggest alternate step.’ Clicking the button would replace the current plan step description with an AI-generated suggestion.
4.1.4. Replanning
At this point, Nóírín’s edits have changed the plan’s trajectory—the next step, which has the agent suggest common themes in papers, is now irrelevant because none of the previous steps retrieve papers. The system automatically detects this and replans subsequent steps accordingly by removing those steps and “autocompleting” the plan with new ones (Figure 6). However, before making these changes, Cocoa notifies Nóírín and asks whether she would like to proceed with replanning; Nóírín accepts. Just like with full plans, Cocoa will present a selection of new steps for Nóírín to choose from. She selects an option to complete the plan.
@@ 4.1.4. Replanning
4.1.5. Adding and deleting steps
In Cocoa, Nóírín can easily add and delete steps from the plan. She notices that the agent has proposed a summary step that may be irrelevant, so she highlights the entire step and hits Backspace to delete it from the plan. This interaction mirrors how one would delete an item in a bulleted list elsewhere in the document. She also thinks it may be useful for the agent to take her ideas from the last step and offer some constructive critiques based on existing literature. She uses the keyboard shortcut Cmd/Ctrl+Enter while selecting the last step to add a new step below it, and edits the step description of the newly added step accordingly. After saving her changes, Nóírín has made all her desired edits to the plan.
@@ 4.2. Co-Execution
Now that Nóírín has made her desired edits to the plan, she is ready to co-execute it with the agent. There are two modes of co-execution: continuous and stepwise. Each step has a step completion indicator that indicates whether the step has not yet been run (a bullet point), is in progress (a spinner), requires user input (a question mark badge), or is complete (a checkmark). Drawing inspiration from computational notebooks, these indicators offer Nóírín a glanceable way to stay informed about the agent’s progress and where she needs to take action in the plan.
@@ 4.2.1. Continuous and stepwise execution
Nóírín wishes to run through the entire plan once to see the output. She triggers continuous co-execution (Figure 7A) with the [Run all] button at the top of the plan. In continuous co-execution, the agent will automatically continue onto the next step as soon as it (or the user) has completed the previous one, until it reaches a user step or the end of the plan.
When she iterates on the plan in subsequent runs to refine specific steps and outputs, Nóírín may opt for stepwise co-execution (Figure 7B). This allows Nóírín to run one plan step at a time, like running individual cells in computational notebooks. When Nóírín hovers her mouse over the bullet point of the first plan step, the bullet point turns into a play button that, when clicked, will run only that step. Unlike some computational notebooks, however, Cocoa does not permit out-of-order stepwise execution [^2] We also note that out-of-order execution in computational notebooks is a major user pain point identified in prior academic work (Chattopadhyay et al., 2020; Lau et al., 2020) and industry reports (Jupyter, 2015; Grus, 2018). because some steps may rely on the outputs of previous ones.
Nóírín can flexibly switch from stepwise to continuous co-execution at any point by clicking the [Run remaining] button at the top of the plan. Conversely, Nóírín can switch from continuous to stepwise co-execution by clicking on the [Pause after this step] button during execution and manually running subsequent steps.
Figure 7. Users can run a plan using continuous execution (left) or stepwise execution (right). A single button click at the top of the plan triggers continuous execution, in which the agent will automatically move onto the next step as soon as it has completed the previous one. To trigger stepwise execution, users hover over the play button beside each step. The agent will not move on to the next step until the user explicitly clicks the play button for that step.
[…]
Continuous and stepwise execution are shown side by side. A single button click at the top of the plan triggers continuous execution, in which the agent will automatically move onto the next step as soon as it has completed the previous one. To trigger stepwise execution, users hover over the play button beside each step. The agent will not move on to the next step until the user explicitly clicks the play button for that step.
@@ 4.2.2. Completing user steps
The agent completes the first step of the plan and arrives at a user step that Nóírín assigned herself. Here, the agent requires Nóírín’s input to continue, and Nóírín can see this because the step has been highlighted in orange. When she clicks into that step, an interactive sidebar opens to the right (Figure 8). The sidebar offers her a built-in paper search functionality connected to the Semantic Scholar academic database, and she uses this to add some relevant papers she has in mind. She clicks Save, adding the papers and their metadata to a plan-specific context pool that the agent references when completing future steps. The agent then proceeds to complete the next step with this updated context to perform a more extensive paper search given the seed papers Nóírín provided.
4.2.3. Editing outputs of agent steps
The agent has now completed the third step, where it conducted a paper search guided by Nóírín’s seed papers from the previous step. Nóírín clicks into that step, opening a sidebar similar to the one she used in the previous step to add papers, this time populated with interactive paper cards. This presents opportunities for her to further guide the agent with her expertise by removing papers she considers irrelevant and adding papers the agent missed. She sees a couple of low-quality results that have incoherent titles and removes them. She then adds one paper she recalls from a past discussion with a collaborator and saves this curated list. The papers in this list, along with their metadata, are stored in the plan’s context pool, accumulated across all completed steps.
The interactive UI in the sidebar dynamically adapts according to the step’s output type, which fall into one of categories in Cocoa: papers, authors, topics, entities, and text. Papers, authors, and topics are drawn from Semantic Scholar and are displayed as interactive cards that Nóírín can add, delete, or search for using the sidebar’s built-in Semantic Scholar search. Entities are lists of items in natural language (search queries, research questions, etc.) and are displayed as editable pills. Text is natural language text displayed in an editable text box.
For any agent step, if the agent fails to return any output, the system will alert Nóírín and ask her to complete the step instead. These features loop in human guidance to bolster the quality of outputs in the face of weak agent performance or agent failures.
[…]
4.2.4. Plan output panel
After the final plan step has been completed, the agent adds a plan output panel (Figure 9) containing a modified version of the last step’s output into the document just below the plan. The modification involves rewriting the output, given the context of the plan and its outputs, to more directly connect to the original user request. This brings the results of running the plan into the document so Nóírín can easily reference it in the context of other content. The panel is collapsed by default to save space and reduce clutter, but she can expand it to copy and paste parts of the output she finds useful and add them to her document. If she does not want the panel in her document at all, she can simply delete it; she can still access the outputs of all steps in the sidebar.

### v4

@@ 2.3. AI for Scientific Research
Significant efforts in recent years have advanced how AI can be used to support scientific research. Of particular interest to us are works that leverage these recent advances to help researchers with planning and execution of literature-focused tasks. Planning a research project is a cognitively demanding, complex process consisting of iterative cycles of divergent and convergent thinking grounded in literature review (Palani et al., 2023; Dean et al., 2006; Chang et al., 2023), making it an ideal scenario to explore human-agent collaboration. One major thread of prior research has focused on leveraging AI models as tools or components in user-driven systems. These include tools for paper reading (Chang et al., 2023; Rachatasumrit et al., 2022) and skimming (Fok et al., 2022), literature review (Lee et al., 2024b; Palani et al., 2023), paper recommendation (Kang et al., 2022b, a), information retrieval and sensemaking (Chan et al., 2018; Kang et al., 2023; Fok et al., 2023; Wang et al., 2024), and ideation (Gu and Krenn, 2024; Baek et al., 2024; Jansen et al., 2024). On the other end of the spectrum are fully automated end-to-end systems where an AI agent has much more agency in planning and executing literature review tasks (Scite, 2025; Elicit, 2024), or even to attempt carrying out entire research projects on their own (Lu et al., 2024; Jansen et al., 2025; Gottweis et al., 2025; Google, 2025; OpenAI, 2025b; Perplexity, 2025; Schmidgall et al., 2025).
@@
As AI systems become more capable, we ask: Is the generation of full research artifacts (e.g., research questions, proposals) the most desired and promising paradigm to assist researchers? Indeed, when working with other (human) collaborators, researchers often find richness in co-evolving partially completed artifacts (Star and Griesemer, 1989; Lee, 2007; Shrum et al., 2007; Jirotka et al., 2013) and answering or asking questions that stimulate critical thinking and reflection (Overholser, 1993; Padesky, 1993). Further, some fully automatically generated artifacts are still found to be of lower quality than those generated with human involvement (Jansen et al., 2025). Schmidgall et al. found that even systems designed for end-to-end research automation produce significantly higher quality outputs with human involvement at each stage (Schmidgall et al., 2025).
@@
In this work, we build on recent advances in AI research support tools, but rather than focusing on generating predetermined research artifacts, we contribute an interaction design pattern that allows for flexible specification of the final artifact and collaboration between a researcher and an AI agent.
@@ 3. Formative Study
Building on prior work, we aim to enhance collaboration between humans and AI agents using a shared, interactive operational representation. We focus on scientific research—a challenging domain for human-agent collaboration given its complex, multi-step workflows and the nuanced decisions researchers make upon processing information-dense scientific papers. Specifically, we investigate the potential for project documents—continuous records of ideas, updates, and tasks—to act as the agent environment. We thus conducted a formative study to answer the following research questions:
@@ 3.1. Participants, Procedure, and Data Analysis
We recruited 9 Ph.D. students (detailed demographics in Table 3 of Appendix A) through an interest form sent to a research organization’s internal Slack channel and the authors’ professional connections. We targeted Ph.D. students as they often lead the detailed planning and execution of research projects and are the primary editors of project documents.
@@
Details for each part can be found in Appendix B. Each participant was given a \$35 USD honorarium after the study. The study was reviewed and exempted by our organization’s internal review board. We qualitatively analyzed project documents produced from the study probe activity alongside study transcripts. Our data analysis methods can be found in Appendix C.
@@ 3.2. Findings
3.2.1. [F1 & F3] Project documents are promising environments for human-agent collaboration
@@ 3.3. Design Goals
    Provide opportunities for human-agent collaboration across planning and execution. Agent developers have called for decoupling planning and execution to avoid fruitless execution (Huyen, 2025). Indeed, participants discussed planning (i.e., breaking down a problem into smaller, more actionable tasks like they did in their project documents) and execution (i.e., completing steps in the plan and refining outputs, with or without AI assistance) as separate but complementary actions. We thus see planning and execution as two distinct stages for fertile researcher-AI collaboration. Because these two stages are closely intertwined and synergistic, we aim to interleave the two in our system.
  • DG2:
    Allow flexible delegation of work between researcher and agent. Researchers may not want to delegate all parts of a plan to AI [F2, F3]. We also observed that this preference may also be context-dependent and constantly shifting. Our system uses a concrete implementation of a theory of flexible delegation proposed by Satyanarayan (Satyanarayan, 2024).
  • DG3:
@@ 4. Cocoa: System Walkthrough and Implementation
We present Cocoa, a system that embeds an AI agent into a document editor and provide an affordance we call interactive plans that allow users to collaboratively plan (co-plan) with the agent—the agent proposes an initial plan of execution to tackle a user request that the user can edit to their liking. Then, users can collaboratively execute (co-execute) the plan with the agent—the user and the agent can build off each other’s work to synergize human and AI capabilities, and provide flexible delegation of human and AI work.
To illustrate the features and functionality of Cocoa, we follow the journey of Nora, a fictional researcher in human-AI interaction, as she uses her project document to further explore open questions in interactive interfaces for AI agents.
@@ 4.1. Co-Planning
Nora’s project document is an informal, reverse chronological log of research progress and includes information such as meeting notes, rough ideas, links to literature, and questions for herself and her collaborators. She has a specific question in her document that she wants to explore further, but being new to the topic, she is not sure how to proceed on her own: “What are new ways AI agents can interactively elicit human feedback?”
4.1.1. Reviewing and selecting an initial plan
To initiate co-planning, she highlights the question to invoke the agent on it using a button that appears in a floating menu. The agent analyzes the question within the context of her entire document and proposes a few different plans via a plan selector UI. Noticing one of the proposed plans started with generating a list of more detailed concepts, which would potentially give her the opportunity to learn and tune the specific paper search directions, Nora selects the plan, which then becomes fully interactive within the document (Figure 3).
@@ 4.1.1. Reviewing and selecting an initial plan
4.1.2. Assigning “Agent steps” and “User steps”
The initial plan included steps for brainstorming relevant topics, searching for papers, and making connections between papers (Figure 3). To balance between effort and control, Nora uses the step assignment toggle (Figure 2 B&C, Figure 3 right) to assign tasks that are relatively low risk but high effort to the agent, such as searching for papers and expanding on some of her preliminary ideas. In contrast, she assigned tasks that were more consequential or required higher-level thinking to herself, such as identifying seed papers for the agent to explore more broadly or making insightful connections between multiple papers.
4.1.3. Plan Editing and Interactive Replanning
Upon further review of the plan, Nora is now also curious about the prominent authors in the area and wish to learn about their papers first. She edits the step description to reflect this. The step description can be edited just like editing any other bulleted list in her document, making this intuitive for Nora. Similarly, Nora can also easily add and delete steps from the plan. For example, she notices that the agent has proposed a summary step that may be irrelevant, so she highlights the step and hits Backspace to delete it. Finally, she uses the keyboard shortcut Cmd/Ctrl+Enter while selecting the last step to add a new step below it, and edits the step description of the newly added step These interactions mirror how one would edit items in a bulleted list in text editors, making them intuitive for Nora. While Cocoa allowed Nora full control over how she wishes to manually edit the plan, this process can be effortful for users. Therefore, the system provides support for replanning in the following ways: First, when editing an individual step, Nora can request the system to provide suggestions for alternatives, informed by the previous steps and document context. Second, after the user has edited some plan step, the plan may become incoherent. For example, a plan step that searches for additional relevant papers given a set of seed papers might no longer make sense if the previous step were edited to produce a list of prominent authors. In this case, Cocoa automatically detects this and replans subsequent steps accordingly by removing those steps and “autocompleting” the plan with new ones (Figure 4).
@@ 4.2. Co-Execution
Now that Nora has made her desired edits to the plan, she is ready to co-execute it with the agent.
@@ 4.2.1. Continuous and stepwise execution
Drawing inspiration from computational notebooks, there are two modes of co-execution: continuous and stepwise. Each step has a step completion indicator that indicates whether the step has not yet been run (a bullet point), is in progress (a spinner), requires user input (a question mark badge), or is complete (a checkmark). This notebook-inspired design allows Nora to dispatch the agent to run autonomously on parts of a plan she trusts the agent to complete on its own. She can also iterate on specific parts of a plan by examining and potentially amending the output of each plan step before continuing.
Nora wishes to run through the entire plan once to see the output. She triggers continuous co-execution with the [Run all] button at the top of the plan. In continuous co-execution, the agent will automatically continue onto the next step as soon as it (or the user) has completed the previous one, until it reaches a user step or the end of the plan. By contrast, if she wishes to iterate and refine the outputs of individual steps before the agent moves onto the next one, she can opt for stepwise co-execution.
Nora can flexibly switch from stepwise to continuous co-execution at any point by clicking the [Run remaining] button at the top of the plan. Conversely, Nora can switch from continuous to stepwise co-execution by clicking on the [Pause after this step] button during execution and manually running subsequent steps. [^2] Cocoa does not permit out-of-order stepwise execution because some steps may rely on the outputs of previous ones. We also note that out-of-order execution in computational notebooks is a major user pain point identified in prior academic work (Chattopadhyay et al., 2020; Lau et al., 2020) and industry reports (Jupyter, 2015; Grus, 2018).
4.2.2. Editing outputs of agent steps
As the agent completed a paper search step, Nora clicks into that step to examine its output. This opens a sidebar to the right that is populated with interactive paper cards (Figure 5). Nora can further guide the agent with her expertise by removing papers she considers irrelevant and/or adding papers the agent missed using a built-in paper search functionality connected to the Semantic Scholar academic database (Kinney et al., 2023). For now, she removes a couple of papers in the output that were not published recently and were published in a less relevant venue.
The interactive UI in the sidebar dynamically adapts according to the step’s output type, which falls into one of five categories in Cocoa: papers, authors, topics, entities, and text. Papers, authors, and topics are drawn from Semantic Scholar and are displayed as interactive cards that Nora can add, delete, or search for using the sidebar’s built-in Semantic Scholar search. Entities are lists of items in natural language (search queries, research questions, etc.) and are displayed as editable pills. Text is natural language text displayed in an editable text box.
If the agent fails to complete its step (e.g., it finds no papers or runs into an error), Cocoa will also proactively alert Nora so that she can complete the step instead, or retry with the agent.
@@ 4.2.3. Completing user steps
Before executing the plan, Nora had assigned step 2 to herself during co-planning. She wanted the agent to start a broad search of many relevant topics and concepts, but then wanted to contribute some seed papers she knew would be relevant to steer the agent’s subsequent actions.
Cocoa has now paused at step 2, highlighting the step in orange to attract Nora’s attention and to seek her input. When she clicks into that step, the same interactive sidebar opens to the right (Figure 5). She could then manually review the topics generated in step 1 and uses the paper search feature built into the sidebar to browse relevant papers and add them to the output of step 2. Additionally, she recalls a few papers that a collaborator had recommended to her previously, finds them using the paper search feature, and adds them to the output.
4.2.4. Final output panel
After the final plan step has been executed, the agent adds a plan output panel (Figure 6) containing a modified version of the last step’s output into the document just below the plan. The modification involves rewriting the output, given the context of the plan and its outputs, to more directly connect to the original user request. This brings the results of running the plan into the document so Nora can easily reference it in the context of other content. The panel is collapsed by default to save space and reduce clutter, but she can expand it to copy and paste parts of the output she finds useful and add them to her document. If she does not want the panel in her document at all, she can simply delete it; she can still access the outputs of all steps in the sidebar.